1. Calculate the electric potential of q1 = 25.0nC at (a) 1.00m, (b) 2.00m, and (c) infinite long distance from it.




2. Calculate the potential energy that another charge q2 = 5.00nC possess when placed at the three different points of the previous example.

1
Expert's answer
2022-02-08T10:28:41-0500

Electric potential="V=k\frac{Q}{r}"

k is Coulomb's constant"9\u00d710^9Nm\u00b2C^{-2}"


a) "V_{1.00m}=\frac{9×10^9×25.0×10^{-9}}{1}=225V"


b)"V_{2.00m}=\frac{9×10^9×25.0×10^{-9}}{2}=112.5V"


c)"V_{inf}=0"


2a) "P.E=\frac{kQ_1Q_2}{r²}"

"P.E_{1m}=\frac{9×10^9×25.0×10^{-9}×5×10^{-9}}{1}=1.125×10^{-6}J"


"P.E_{2}=\frac{9×10^9×25.0×10^{-9}×5×10^{-9}}{2^2}=2.8125×10^{-7}J"



c)"P.E_{inf}=0"